Hello,
I have a word problem on travel and distance.
Two cars are 400 miles apart. Both start at the same time and travel toward one another. They meet 4 hours later. If the speed of one car is 20 miles per hour faster than the other, what is the speed of each car?
Thank you

Answer by rfer(12678)

x+x+20=400/4
2x+20=100
2x=80
x=40 mph
x+20=60 mph